# Rethinking Graph Structure Learning in the Era of LLMs

### v4

@@ 4.3 Leaf-oriented Two-step Sampling
Motivation. Although structural optimization can be applied to all leaf nodes, we adopt a two-step sampling technology to balance running efficiency and practical performance. Specifically, we first identify a limited set of nodes requiring optimization from a topological perspective. Subsequently, we select a candidate set of nodes closely related to these nodes from a text semantic perspective. Finally, through edge addition or removal, we facilitate training-free GSL. The detailed algorithm of the two-step sampling can ben found in Appendix E.3.
@@
$P_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}},\;\;P_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}},\gamma\neq\alpha}{\sum}}\!\!\!{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha}))}}.$ (7)
@@ 5.2 Performance Comparison
Node Clustering. To further answer Q1, we conducted a comprehensive comparison of LLaTA with GSL methods that have demonstrated strong performance in node classification, extending to node clustering. As shown in Table 3, LLaTA consistently outperforms competing methods across all four datasets, significantly enhancing the effectiveness of GNNs in unsupervised tasks. These results confirm that LLaTA is applicable to diverse tasks, demonstrating its generalization and adaptability.
Table 1: Node classification accuracy(%) on 11 TAG datasets. Highlighted are the top first, second, and third accuracy. "OOM" denotes out of memory and "OOT" denotes out of time.
@@ 5.4 Robustness Analysis
Table 4: Improvements of LLaTA over different GNN/LLM-GNN backbones.
@@
Table 5: Time complexity analysis of LLaTA and other GSL method.
@@ 6 Conclusion
In this paper, we first introduce a novel optimization framework for GSL by rethinking its integration with LLMs, addressing the challenges of incorporating textual information. Subsequently, We propose LLaTA as an instantiation of this novel framework, which leverages a structural encoding tree to achieve efficient LLM in-context learning. This enables the LLM to comprehensively understand both topology and text insights from the original graph, ultimately relying on reliable inference to obtain improved graph structure. To this end, LLaTA eliminates the need for costly fine-tuning and achieves SOTA performance. Inspired by the noise experiments, a promising direction is developing more robust tree optimization algorithms, laying a solid foundation for LLM-empowered GSL.
@@ Appendix A Encoding Tree and Structural Entropy
where $g_{\alpha}$ represents the sum of the weights of cross-node edges that connect nodes within partition $\mathcal{C}_{\alpha}$ to nodes outside $\mathcal{C}_{\alpha}$. $\operatorname{vol}(\alpha)$ denotes the sum of the degrees of all nodes within $\mathcal{C}_{\alpha}$. $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ is the structural entropy of node $\alpha$, representing the $K$-dimensional structural entropy of $\alpha$ when the height of $\mathcal{T}$ is $K$. The core of this measurement lies in the observation that, in a highly connected graph, nodes frequently interact with their neighbors. By employing random walks, these interactions can be captured, and entropy can be introduced as a measure of topology uncertainty [12]. Specifically, the one-dimensional structural measurement of $\mathcal{G}$ can be quantified using the stationary distribution of its degrees $d$ and Shannon entropy. This concept can be generalized to $K$-dimensional measurements using the structural encoding tree in Definition 1.
@@ Appendix B Analysis of Hierarchical Community
While low-level communities are primarily focused on capturing local topological structure, they are not simple local neighborhoods. Under the global structural entropy minimization principle[12], the construction of the encoding tree ensures that each module, including those at the bottom level, is formed in a way that minimizes global uncertainty and boundary complexity.
From the random walk perspective, the structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ quantifies the average uncertainty of a walker’s position in the network under a hierarchical modular encoding. As shown in [12], the structural entropy with respect to a flat partition $\mathcal{P}$ can be decomposed as:
@@ C.1 Existing GSL Paradigms
Coupled Paradigm. The graph learner and backbone are tightly coupled [2, 6, 28] to learn improved structure by specific downstream tasks (e.g., node, link, and graph level). Limitations: (1) Unstable Performance: This task- and backbone-specific architecture limits the generalizability of the improved structure. Moreover, the strong dependency between the structure learning module and the downstream GNN makes performance vulnerable to low-quality graphs: if the original graph is poor, the resulting low-quality embeddings from the GNN directly degrade the effectiveness of the structure learning module. (2) Inflexible Deployment: Switching the downstream backbone or task requires retraining, which significantly limits its adaptability and scalability.
Decoupled Paradigm. The graph learner and downstream backbone are trained independently [17, 15, 1], avoiding co-training, thereby enabling better compatibility with LLMs for GSL. We observe that recent LLM-based GSL methods also tend to adopt this decoupled paradigm [7, 34]. However, these methods exhibit a different form of coupling (i.e., fine-tuning based on the pre-defined instruction datasets), which leads to the following issues: (1) They fail to directly and jointly incorporate topology and text, relying instead on instruction datasets. Moreover, they are highly sensitive to the quality of these datasets. (2) They rely on complex mechanisms and significant time to construct instruction datasets and fine-tuning. Therefore, from a model design perspective, we recommend adopting a comprehensive decoupled paradigm to integrate GSL and LLMs, minimizing reliance on fine-tuning and instead prioritizing more efficient and reliable inference with high-quality in-context prompts.
@@ C.2 Comparative Analysis with SE-GSL and LLM-based Methods
In this section, we provide a detailed comparative analysis between our proposed LLaTA and existing approaches, including SE-GSL [42], GraphEdit [7], LLM4RGNN [34], and LangGSL [24]. The goal is to highlight the fundamental differences in design and demonstrate the advantages of our method.
@@ Appendix D Emperical Study
Over-smoothing and Over-squashing. They are two common challenges in graph learning [20, 38]. Specifically, over-smoothing occurs when node features become indistinguishable as they converge to similar values, while over-squashing arises when nodes fail to capture information from distant neighbors, particularly in the presence of local bottlenecks in the graph. Both issues are strongly tied to the quality of the graph structure. To evaluate graph structure quality in light of these challenges, we employ the following analysis method: To begin with, we train a GCN on the improved graph to generate node embeddings. Then, we compute over-smoothing and over-squashing values by randomly sampling several node pairs: (1) Over-smoothing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of heterophilous nodes and compute their average cosine similarity based on the GCN-generated node embeddings. (2) Over-squashing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of distant homophilous nodes and compute their average cosine similarity using the same embeddings.
@@ E.1 Structural Entropy Minimization
1Input: A graph $\mathcal{G}$, the height of encoding tree $K$.
2 Output: Encoding tree $\mathcal{T}^{K}$ with height $K$.
3 // Initialize encoding tree $\mathcal{T}$ with height 1
4 Create root node $\lambda$;
5 for $v_{i}\in\mathcal{V}$ do
6 $\mathrm{NI}_{\mathcal{T}}(v_{i},\alpha_{i})$ according to Definition. 3;
7// Node combining
8 while $\lambda$ has more than 2 children do
9   Select $\alpha$ and $\beta$ in $\mathcal{T}$, conditioned on $\alpha^{+}=\beta^{+}=\lambda$ and $\underset{\alpha,\beta}{\arg\max}\left(\mathcal{H}^{\mathcal{T}}(\mathcal{G})-\mathcal{H}_{\mathrm{NC}_{\mathcal{T}}(\alpha,\beta)}^{\mathcal{T}}(\mathcal{G})\right)$;
10 $\mathrm{NC}_{\mathcal{T}}(\alpha,\beta)$ according to Definition. 4;
11// Node lifting
@@ 4 Create root node $\lambda$;
13   Select non-root nodes $\alpha$ and $\beta$ in $\mathcal{T}$, conditioned on $\alpha^{+}=\beta$ and
14 $\underset{\alpha,\beta}{\arg\max}\left(\mathcal{H}^{\mathcal{T}}(\mathcal{G})-\mathcal{H}_{\mathrm{NL}_{\mathcal{T}}(\alpha,\beta)}^{\mathcal{T}}(\mathcal{G})\right)$;
15 $\mathrm{NL}_{\mathcal{T}}(\alpha,\beta)$ according to Definition. 5;
16
17Return $\mathcal{T}^{K}\leftarrow\mathcal{T}$;
@@ E.2 Tree-based Adaptive Clustering
To optimize the structural encoding tree obtained by Appendix E.1 using text-driven node features, we introduce an adaptive clustering approach based on tree structures and the silhouette coefficient. This method reallocates leaf nodes according to text-driven insights from LLMs. By maximizing homophily within communities, we ensure that nodes with similar label classes are effectively grouped, aligning with structural patterns commonly observed in real-world applications. For graphs exhibiting heterophily, which are less common but still present, recent studies have demonstrated that uncovering their intrinsic higher-order homophilous patterns is an effective strategy [4, 5, 23].
@@
1Input: Encoding tree $\mathcal{T}$, soft labels $\mathbf{Y}^{cls}$ for each leaf node, hyperparameter $s$.
2 Output: Optimized encoding tree $\mathcal{T}^{\star}$.
3 for $\mathcal{C}^{\ell}\in\mathcal{T}$ do
4   Let $\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}=\{y^{cls}_{1},y^{cls}_{2},\dots,y^{cls}_{m}\}$ be the soft labels of all nodes in $\mathcal{C}^{\ell}$;
5   Let $k^{\star}=0$ and $sil_{max}=-1$;
6 for $2<k<\frac{m}{2}$ do
7      Perform k-means($\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}$, $k$) and calculate average silhouette coefficient $sil_{k}$ of all clusters;
8    if $sil_{k}-sil_{max}<s$ then
9       break;
10    else
11       $sil_{max}=sil_{k}$; $k^{\star}=k$;
12
13  Perform k-means($\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}$, $k^{\star}$) and get a set of clusters $\{C_{1},C_{2},\dots,C_{k^{\star}}\}$;
14 for $C_{i}\in\{C_{1},C_{2},\dots,C_{k^{\star}}\}$ do
15    if $|C_{i}|>1$ then
16         Construct the leaf nodes in $C_{i}$ as a new community and set the parent node of the new community to be the same as the parent node of $\mathcal{C}^{\ell}$;
17    else
18         Reallocate the leaf node in $C_{i}$ to other low-level communities based on $\mathbf{Y}^{cls}$;
19
20
21
22Return $\mathcal{T}^{\star}\leftarrow\mathcal{T}$;
@@ E.3 Leaf-oriented Two-step Sampling
1Input: Original Graph $\mathcal{G}$, Optimized encoding tree $\mathcal{T}^{\star}$, soft label $\mathbf{Y}^{cls}$ for each leaf node, hyperparameters $\theta$, $r$.
2 Output: Optimized graph $\mathcal{G}^{\star}$.
3for $\mathcal{C}^{\ell}\in\mathcal{T}^{\star}$ do
4 for $i=1$ to $m\times r$ do
5      // Step 1: Sample a leaf node $\alpha$ from $\mathcal{C}^{\ell}$
6      Let $set_{1}=\{\alpha|\alpha\in\mathcal{C}^{\ell}\}$;
7    for $\alpha\in set_{1}$ do
8         Compute $\mathcal{H}^{\mathcal{T}^{\star}}(\mathcal{G},\alpha)$ according to Eq. (9);
9         Compute $P_{topo}(\alpha)$ according to Eq. (7);
10     Sample a leaf node $\alpha$ based on $P_{topo}(\alpha)$ from $set_{1}$;
11      // Step 2: Sample an edge for $\alpha$
12      Let $set_{2}=\{\beta|\beta\in\mathcal{C}^{\ell}$ and $\beta\neq\alpha\}$;
13    if $|set_{2}|<\theta$ then
14         // Expand the candidate set
15         Let $\phi$ be the grandparent of $\alpha$;
16       for $\gamma\in\phi$ and $\gamma\neq\alpha^{+}$ do
17            Compute $y_{\gamma}^{cls}=\frac{1}{m}\sum_{\alpha\in\gamma}y_{\alpha}^{cls}$;
18            Compute $\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha})$ according to Eq. (3);
19       for $|set_{2}|<\theta$ do
20            Choose low-level community with highest $\mathrm{sim}$ and add its children to $set_{2}$;
21
22
23    for $\beta\in set_{2}$ do
24         Compute $\mathrm{sim}(y^{cls}_{\beta},y^{cls}_{\alpha})$ according to Eq. (3);
25         Compute $P_{sema}^{\alpha}(\beta)$ according to Eq. (7);
26
27
28    if $|set_{2}|>\theta$ then
29         Keep the top $\theta$ nodes in $set_{2}$ based on the soft label similarity with $\alpha$.
30
31     For edge addition, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in descending order;
32      For edge removal, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in ascending order;
33      Add edge to or remove edge from graph $\mathcal{G}$;
34
35
36Return $\mathcal{G}^{\star}\leftarrow\mathcal{G}$;

### v5

@@ 4.3 Leaf-oriented Two-step Sampling
Motivation. Although structural optimization could be applied to all leaf nodes, we adopt a two-step sampling strategy to balance efficiency and performance. First, we select a small set of nodes that require optimization from a topological perspective. Second, we sample textually related candidate nodes for them. We then update the graph via edge additions or removals in a training-free manner. The full two-step sampling algorithm is given in Appendix E.3.
@@
$P_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}},$ (9)
$P_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}},\gamma\neq\alpha}{\sum}}\!\!\!{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha}))}}.$ (10)
@@ 5.2 Performance Comparison
Node Clustering. To further answer Q1, we compare LLaTA with strong GSL baselines for node classification, extended to node clustering. As shown in Table 2, LLaTA consistently outperforms all competitors on four datasets, markedly improving GNNs on unsupervised tasks. These results confirm that LLaTA is applicable to diverse tasks, demonstrating its generalization and adaptability.
Table 1: Node classification accuracy(%) on 11 TAG datasets. Highlighted are the top first, second, and third accuracy. ”OOM” denotes out of memory and ”OOT” denotes out of time.
@@ 5.4 Robustness Analysis
Table 4: Improvements of LLaTA over different backbones.
@@
Table 5: Time complexity analysis of LLaTA and baselines.
@@ 6 Conclusion
In this paper, we propose a tree-based optimization framework for GSL that explicitly integrates LLMs and textual information. We instantiate this framework as LLaTA, which leverages a structural encoding tree for efficient LLM in-context learning, enabling joint understanding of topology and text and yielding improved graph structures without costly fine-tuning. LLaTA achieves SOTA performance across multiple benchmarks. Inspired by the noise experiments, future work will explore more robust tree optimization algorithms for LLM-empowered GSL.
@@ Appendix A Encoding Tree and Structural Entropy
where $g_{\alpha}$ represents the sum of the weights of cross-node edges that connect nodes within partition $\mathcal{C}_{\alpha}$ to nodes outside $\mathcal{C}_{\alpha}$. $\operatorname{vol}(\alpha)$ denotes the sum of the degrees of all nodes within $\mathcal{C}_{\alpha}$. $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ is the structural entropy of node $\alpha$, representing the $K$-dimensional structural entropy of $\alpha$ when the height of $\mathcal{T}$ is $K$. The core of this measurement lies in the observation that, in a highly connected graph, nodes frequently interact with their neighbors. By employing random walks, these interactions can be captured, and entropy can be introduced as a measure of topology uncertainty (Li and Pan, 2016). Specifically, the one-dimensional structural measurement of $\mathcal{G}$ can be quantified using the stationary distribution of its degrees $d$ and Shannon entropy. This concept can be generalized to $K$-dimensional measurements using the structural encoding tree in Definition A.1.
@@ Appendix B Analysis of Hierarchical Community
While low-level communities are primarily focused on capturing local topological structure, they are not simple local neighborhoods. Under the global structural entropy minimization principle(Li and Pan, 2016), the construction of the encoding tree ensures that each module, including those at the bottom level, is formed in a way that minimizes global uncertainty and boundary complexity.
From the random walk perspective, the structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G})$ quantifies the average uncertainty of a walker’s position in the network under a hierarchical modular encoding. As shown in (Li and Pan, 2016), the structural entropy with respect to a flat partition $\mathcal{P}$ can be decomposed as:
@@ C.1 Existing GSL Paradigms
Coupled Paradigm. The graph learner and backbone are tightly coupled (Chen et al., 2020; Fatemi et al., 2021; Wu et al., 2023) to learn improved structure by specific downstream tasks (e.g., node, link, and graph level). Limitations: (1) Unstable Performance: This task- and backbone-specific architecture limits the generalizability of the improved structure. Moreover, the strong dependency between the structure learning module and the downstream GNN makes performance vulnerable to low-quality graphs: if the original graph is poor, the resulting low-quality embeddings from the GNN directly degrade the effectiveness of the structure learning module. (2) Inflexible Deployment: Switching the downstream backbone or task requires retraining, which significantly limits its adaptability and scalability.
Decoupled Paradigm. The graph learner and downstream backbone are trained independently (Liu et al., 2022b, a; Bi et al., 2024), avoiding co-training, thereby enabling better compatibility with LLMs for GSL. We observe that recent LLM-based GSL methods also tend to adopt this decoupled paradigm (Guo et al., 2024; Zhang et al., 2024). However, these methods exhibit a different form of coupling (i.e., fine-tuning based on the pre-defined instruction datasets), which leads to the following issues: (1) They fail to directly and jointly incorporate topology and text, relying instead on instruction datasets. Moreover, they are highly sensitive to the quality of these datasets. (2) They rely on complex mechanisms and significant time to construct instruction datasets and fine-tuning. Therefore, from a model design perspective, we recommend adopting a comprehensive decoupled paradigm to integrate GSL and LLMs, minimizing reliance on fine-tuning and instead prioritizing more efficient and reliable inference with high-quality in-context prompts.
@@ C.2 Comparative Analysis with SE-GSL and LLM-based Methods
In this section, we provide a detailed comparative analysis between our proposed LLaTA and existing approaches, including SE-GSL (Zou et al., 2023), GraphEdit (Guo et al., 2024), LLM4RGNN (Zhang et al., 2024), and LangGSL (Su et al., 2024). The goal is to highlight the fundamental differences in design and demonstrate the advantages of our method.
@@ Appendix D Emperical Study
Over-smoothing and Over-squashing. They are two common challenges in graph learning (Nguyen et al., 2023; Zheng et al., 2024). Specifically, over-smoothing occurs when node features become indistinguishable as they converge to similar values, while over-squashing arises when nodes fail to capture information from distant neighbors, particularly in the presence of local bottlenecks in the graph. Both issues are strongly tied to the quality of the graph structure. To evaluate graph structure quality in light of these challenges, we employ the following analysis method: To begin with, we train a GCN on the improved graph to generate node embeddings. Then, we compute over-smoothing and over-squashing values by randomly sampling several node pairs: (1) Over-smoothing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of heterophilous nodes and compute their average cosine similarity based on the GCN-generated node embeddings. (2) Over-squashing: We randomly sample $0.2\times|\mathcal{V}|$ pairs of distant homophilous nodes and compute their average cosine similarity using the same embeddings.
@@ E.1 Structural Entropy Minimization
Input: A graph $\mathcal{G}$, the height of encoding tree $K$.
Output: Encoding tree $\mathcal{T}^{K}$ with height $K$.
// Initialize encoding tree $\mathcal{T}$ with height 1
Create root node $\lambda$;
for $v_{i}\in\mathcal{V}$ do
 $\mathrm{NI}_{\mathcal{T}}(v_{i},\alpha_{i})$ according to Definition. E.1;
// Node combining
while $\lambda$ has more than 2 children do
   Select $\alpha$ and $\beta$ in $\mathcal{T}$, conditioned on $\alpha^{+}=\beta^{+}=\lambda$ and $\underset{\alpha,\beta}{\arg\max}\left(\mathcal{H}^{\mathcal{T}}(\mathcal{G})-\mathcal{H}_{\mathrm{NC}_{\mathcal{T}}(\alpha,\beta)}^{\mathcal{T}}(\mathcal{G})\right)$;
$\mathrm{NC}_{\mathcal{T}}(\alpha,\beta)$ according to Definition. E.2;
// Node lifting
@@
   Select non-root nodes $\alpha$ and $\beta$ in $\mathcal{T}$, conditioned on $\alpha^{+}=\beta$ and
$\underset{\alpha,\beta}{\arg\max}\left(\mathcal{H}^{\mathcal{T}}(\mathcal{G})-\mathcal{H}_{\mathrm{NL}_{\mathcal{T}}(\alpha,\beta)}^{\mathcal{T}}(\mathcal{G})\right)$;
$\mathrm{NL}_{\mathcal{T}}(\alpha,\beta)$ according to Definition. E.3;
Return $\mathcal{T}^{K}\leftarrow\mathcal{T}$;
@@ E.2 Tree-based Adaptive Clustering
To optimize the structural encoding tree obtained by Appendix E.1 using text-driven node features, we introduce an adaptive clustering approach based on tree structures and the silhouette coefficient. This method reallocates leaf nodes according to text-driven insights from LLMs. By maximizing homophily within communities, we ensure that nodes with similar label classes are effectively grouped, aligning with structural patterns commonly observed in real-world applications. For graphs exhibiting heterophily, which are less common but still present, recent studies have demonstrated that uncovering their intrinsic higher-order homophilous patterns is an effective strategy (Dai et al., 2022; Du et al., 2022; Song et al., 2023).
@@
Input: Encoding tree $\mathcal{T}$, soft labels $\mathbf{Y}^{cls}$ for each leaf node, hyperparameter $s$.
Output: Optimized encoding tree $\mathcal{T}^{\star}$.
for $\mathcal{C}^{\ell}\in\mathcal{T}$ do
   Let $\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}=\{y^{cls}_{1},y^{cls}_{2},\dots,y^{cls}_{m}\}$ be the soft labels of all nodes in $\mathcal{C}^{\ell}$;
Let $k^{\star}=0$ and $sil_{max}=-1$;
for $2<k<\frac{m}{2}$ do
      Perform k-means($\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}$, $k$) and calculate average silhouette coefficient $sil_{k}$ of all clusters;
if $sil_{k}-sil_{max}<s$ then
       break;
    else
       $sil_{max}=sil_{k}$; $k^{\star}=k$;
  Perform k-means($\mathbf{Y}^{cls}_{\mathcal{C}^{\ell}}$, $k^{\star}$) and get a set of clusters $\{C_{1},C_{2},\dots,C_{k^{\star}}\}$;
for $C_{i}\in\{C_{1},C_{2},\dots,C_{k^{\star}}\}$ do
    if $|C_{i}|>1$ then
         Construct the leaf nodes in $C_{i}$ as a new community and set the parent node of the new community to be the same as the parent node of $\mathcal{C}^{\ell}$;
    else
         Reallocate the leaf node in $C_{i}$ to other low-level communities based on $\mathbf{Y}^{cls}$;
Return $\mathcal{T}^{\star}\leftarrow\mathcal{T}$;
@@ E.3 Leaf-oriented Two-step Sampling
Input: Original Graph $\mathcal{G}$, Optimized encoding tree $\mathcal{T}^{\star}$, soft label $\mathbf{Y}^{cls}$ for each leaf node, hyperparameters $\theta$, $r$.
Output: Optimized graph $\mathcal{G}^{\star}$.
for $\mathcal{C}^{\ell}\in\mathcal{T}^{\star}$ do
 for $i=1$ to $m\times r$ do
      // Step 1: Sample a leaf node $\alpha$ from $\mathcal{C}^{\ell}$
Let $set_{1}=\{\alpha|\alpha\in\mathcal{C}^{\ell}\}$;
for $\alpha\in set_{1}$ do
         Compute $\mathcal{H}^{\mathcal{T}^{\star}}(\mathcal{G},\alpha)$ according to Eq. (12);
Compute $P_{topo}(\alpha)$ according to Eq. (9);
     Sample a leaf node $\alpha$ based on $P_{topo}(\alpha)$ from $set_{1}$;
// Step 2: Sample an edge for $\alpha$
Let $set_{2}=\{\beta|\beta\in\mathcal{C}^{\ell}$ and $\beta\neq\alpha\}$;
if $|set_{2}|<\theta$ then
         // Expand the candidate set
Let $\phi$ be the grandparent of $\alpha$;
for $\gamma\in\phi$ and $\gamma\neq\alpha^{+}$ do
            Compute $y_{\gamma}^{cls}=\frac{1}{m}\sum_{\alpha\in\gamma}y_{\alpha}^{cls}$;
Compute $\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha})$;
       for $|set_{2}|<\theta$ do
            Choose low-level community with highest $\mathrm{sim}$ and add its children to $set_{2}$;
    for $\beta\in set_{2}$ do
         Compute $\mathrm{sim}(y^{cls}_{\beta},y^{cls}_{\alpha})$;
Compute $P_{sema}^{\alpha}(\beta)$ according to Eq. (10);
    if $|set_{2}|>\theta$ then
         Keep the top $\theta$ nodes in $set_{2}$ based on the soft label similarity with $\alpha$.
     For edge addition, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in descending order;
For edge removal, sample a leaf node $\beta$ from $set_{2}$ based on $P_{sema}^{\alpha}(\beta)$ ranked in ascending order;
Add edge to or remove edge from graph $\mathcal{G}$;
Return $\mathcal{G}^{\star}\leftarrow\mathcal{G}$;